please help me with this.I do not have any idea how to begin solving this.
f(x)=x^2-2x+3, g(x)=-2x+2 and h(x)=|x-3| What is the correct evaluation of f(2)?
thank you

Answer by stanbon(75887)

f(x)=x^2-2x+3, g(x)=-2x+2 and h(x)=|x-3|
What is the correct evaluation of f(2)?
----------
f(x) = x^2-2x+3
f(2) = 2^2-2*2+3
f(2) = 4-4+3
f(2) = 3
===============
Cheers,
Stan H.
===============